### v1

# DreamGaussian: Generative Gaussian Splatting for Efficient 3D Content Creation

Abstract
Recent advances in 3D content creation mostly leverage optimization-based 3D generation via score distillation sampling (SDS). Though promising results have been exhibited, these methods often suffer from slow per-sample optimization, limiting their practical usage. In this paper, we propose DreamGaussian, a novel 3D content generation framework that achieves both efficiency and quality simultaneously. Our key insight is to design a generative 3D Gaussian Splatting model with companioned mesh extraction and texture refinement in UV space. In contrast to the occupancy pruning used in Neural Radiance Fields, we demonstrate that the progressive densification of 3D Gaussians converges significantly faster for 3D generative tasks. To further enhance the texture quality and facilitate downstream applications, we introduce an efficient algorithm to convert 3D Gaussians into textured meshes and apply a fine-tuning stage to refine the details. Extensive experiments demonstrate the superior efficiency and competitive generation quality of our proposed approach. Notably, DreamGaussian produces high-quality textured meshes in just 2 minutes from a single-view image, achieving approximately 10 times acceleration compared to existing methods.
Figure 1: DreamGaussian aims at accelerating the optimization process of both image- and text-to-3D tasks. We are able to generate a high quality textured mesh in several minutes.
1 Introduction
Automatic 3D digital content creation finds applications across various domains, including digital games, advertising, films, and the MetaVerse. The core techniques, including image-to-3D and text-to-3D, offer substantial advantages by significantly reducing the need for manual labor among professional artists and empowering non-professional users to engage in 3D asset creation. Drawing inspiration from recent breakthroughs in 2D content generation (Rombach et al., 2022), the field of 3D content creation has experienced rapid advancements. Recent studies in 3D creation can be classified into two principal categories: inference-only 3D native methods and optimization-based 2D lifting methods. Theoretically, 3D native methods (Jun & Nichol, 2023; Nichol et al., 2022; Gupta et al., 2023) exhibit the potential to generate 3D-consistent assets within seconds, albeit at the cost of requiring extensive training on large-scale 3D datasets. The creation of such datasets demand substantial human effort, and even with these efforts, they continue to grapple with issues related to limited diversity and realism (Deitke et al., 2023b; a; Wu et al., 2023).
On the other hand, Dreamfusion (Poole et al., 2022) proposes Score Distillation Sampling (SDS) to address the 3D data limitation by distilling 3D geometry and appearance from powerful 2D diffusion models (Saharia et al., 2022), which inspires the development of recent 2D lifting methods (Lin et al., 2023; Wang et al., 2023b; Chen et al., 2023b). In order to cope with the inconsistency and ambiguity caused by the SDS supervision, Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) are usually adopted for their capability in modeling rich 3D information. Although the generation quality has been increasingly improved, these approaches are notorious for hours-long optimization time due to the costly NeRF rendering, which restricts them from being deployed to real-world applications at scale. We argue that the occupancy pruning technique used to accelerate NeRF (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022) is ineffective in generative settings when supervised by the ambiguous SDS loss as opposed to reconstruction settings.
In this work, we introduce the DreamGaussian framework, which greatly improves the 3D content generation efficiency by refining the design choices in an optimization-based pipeline. Photo-realistic 3D assets with explicit mesh and texture maps can be generated from a single-view image within only 2 minutes using our method. Our core design is to adapt 3D Gaussian Splatting (Kerbl et al., 2023) into the generative setting with companioned meshes extraction and texture refinement. Compared to previous methods with the NeRF representation, which find difficulties in effectively pruning empty space, our generative Gaussian splatting significantly simplifies the optimization landscape. Specifically, we demonstrate the progressive densification of Gaussian splitting, which highly in accordance with the optimization progress of generative settings, greatly improves the generation efficiency. As illustrated in Figure 1, our image-to-3D pipeline swiftly produces a coarse shape within seconds and converges efficiently in around $500$ steps on a single GPU.
Due to the ambiguity in SDS supervision and spatial densification, the directly generated results from 3D Gaussians tend to be blurry. To address the issue, we identify that the texture needs to be refined explicitly, which requires delicate textured polygonal mesh extraction from the generated 3D Gaussians. While this task has not been explored before, we design an efficient algorithm for mesh extraction from 3D Gaussians by local density querying. Then a generative UV-space refinement stage is proposed to enhance the texture details. Given the observation that directly applying the latent space SDS loss as in the first stage results in over-saturated blocky artifacts on the UV map, we take the inspiration from diffusion-based image editing methods (Meng et al., 2021) and perform image space supervision. Compared to previous texture refinement approaches, our refinement stage achieves better fidelity while keeping high efficiency.
In summary, our contributions are:
  1. 1.
    We propose a novel framework for 3D content creation by adapting Gaussian splatting into generative settings, significantly reducing the generation time of optimization-based 2D lifting methods.
  2. 2.
    We design an efficient mesh extraction algorithm from 3D Gaussians and a UV-space texture refinement stage to further enhance the generation quality.
  3. 3.
    Extensive experiments on both Image-to-3D and Text-to-3D tasks demonstrate that our method effectively balances optimization time and generation fidelity, unlocking new possibilities for real-world deployment of 3D content generation.
2 Related Work
2.1 3D Representations
Various 3D representations have been proposed for different 3D tasks. Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) employs a volumetric rendering and has been popular for enabling 3D optimization with only 2D supervision. Although NeRF has become widely used in both 3D reconstruction (Barron et al., 2022; Li et al., 2023c; Chen et al., 2022; Hedman et al., 2021) and generation (Poole et al., 2022; Lin et al., 2023; Chan et al., 2022), optimizing NeRF can be time-consuming. Various attempts have been made to accelerate the training of NeRF (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022), but these works only focus on the reconstruction setting. The common technique of spatial pruning fails to accelerate the generation setting. Recently, 3D Gaussian splatting (Kerbl et al., 2023) has been proposed as an alternative 3D representation to NeRF, which has demonstrated impressive quality and speed in 3D reconstruction (Luiten et al., 2023). The efficient differentiable rendering implementation and model design enables fast training without relying on spatial pruning. In this work, we for the first time adapt 3D Gaussian splatting into generation tasks to release the potential of optimization-based methods.
2.2 Text-to-3D Generation
Text-to-3D generation aims at generating 3D assets from a text prompt. Recently, data-driven 2D diffusion models have achieved notable success in text-to-image generation (Ho et al., 2020; Rombach et al., 2022; Saharia et al., 2022). However, transferring it to 3D generation is non-trivial due to the challenge of curating large-scale 3D datasets. Existing 3D native diffusion models usually work on a single object category and suffer from limited diversity (Jun & Nichol, 2023; Nichol et al., 2022; Gupta et al., 2023; Lorraine et al., 2023; Zhang et al., 2023; Zheng et al., 2023; Ntavelis et al., 2023; Chen et al., 2023a; Cheng et al., 2023; Gao et al., 2022). To achieve open-vocabulary 3D generation, severl methods propose to lift 2D image models for 3D generation (Jain et al., 2022; Poole et al., 2022; Wang et al., 2023a; Mohammad Khalid et al., 2022; Michel et al., 2022). Such 2D lifting methods optimize a 3D representation to achieve a high likelihood in pretrained 2D diffusion models when rendered from different viewpoints, such that both 3D consistency and realisticity can be ensured. Following works continue to enhance various aspects such as generation fidelity and training stability (Lin et al., 2023; Tsalicoglou et al., 2023; Zhu & Zhuang, 2023; Yu et al., 2023; Li et al., 2023b; Chen et al., 2023c; Wang et al., 2023b; Huang et al., 2023; Metzer et al., 2022; Chen et al., 2023b), and explore further applications (Zhuang et al., 2023; Singer et al., 2023; Raj et al., 2023). However, these optimization-based 2D lifting approaches usually suffer from long per-case optimization time. Particularly, employing NeRF as the 3D representation leads to expensive computations during both forward and backward. In this work, we choose 3D Gaussians as the differentiable 3D representation and empirically show that it has a simpler optimization landscape.
2.3 Image-to-3D Generation
Image-to-3D generation targets generating 3D assets from a reference image. The problem can also be formulated as single-view 3D reconstruction (Yu et al., 2021; Trevithick & Yang, 2021; Duggal & Pathak, 2022), but such reconstruction settings usually produce blurry results due to the lack of uncertainty modeling. Text-to-3D methods can also be adapted for image-to-3D generation (Xu et al., 2023a; Tang et al., 2023b; Melas-Kyriazi et al., 2023) using image captioning models (Li et al., 2022; 2023a). Recently, Zero-1-to-3 (Liu et al., 2023b) explicitly models the camera transformation into 2D diffusion models and enable zero-shot image-conditioned novel view synthesis. It achieves high 3D generation quality when combined with SDS, but still suffers from long optimization time (Tang, 2022; Qian et al., 2023). One-2-3-45 (Liu et al., 2023a) trains a multi-view reconstruction model for acceleration at the cost of the generation quality. With an efficiency-optimized framework, our work shortens the image-to-3D optimization time to 2 minutes with little sacrifice on quality.
3 Our Approach
In this section, we introduce our two-stage framework for efficient 3D content generation for both Image-to-3D and Text-to-3D tasks as illustrated in Figure 2. Firstly, we adapt 3D Gaussian splatting (Kerbl et al., 2023) into generation tasks for efficient initialization through SDS (Poole et al., 2022) (Section 3.1). Next, we propose an efficient algorithm to extract a textured mesh from 3D Gaussians (Section 3.2). This texture is then fine-tuned by differentiable rendering (Laine et al., 2020) through a UV-space refinement stage (Section 3.3) for final exportation.
Figure 2: DreamGaussian Framework. 3D Gaussians are used for efficient initialization of geometry and appearance using single-step SDS loss. We then extract a textured mesh and refine the texture image with a multi-step MSE loss.
3.1 Generative Gaussian Splatting
Gaussian splatting (Kerbl et al., 2023) represents 3D information with a set of 3D Gaussians. It has been proven effective in reconstruction settings (Kerbl et al., 2023; Luiten et al., 2023) with high inference speed and reconstruction quality under similar modeling time with NeRF. However, its usage in a generative manner has not been explored. We identify that the 3D Gaussians can be efficient for 3D generation tasks.
Specifically, the location of each Gaussian can be described with a center $\mathbf{x}\in\mathbb{R}^{3}$, a scaling factor $\mathbf{s}\in\mathbb{R}^{3}$, and a rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$. We also store an opacity value $\alpha\in\mathbb{R}$ and a color feature $\mathbf{c}\in\mathbb{R}^{3}$ for volumetric rendering. Spherical harmonics are disabled since we only want to model simple diffuse color. All the above optimizable parameters is presented by ${\Theta}$, where ${\Theta}_{i}=\{\mathbf{x}_{i},\mathbf{s}_{i},\mathbf{q}_{i},\alpha_{i},\mathbf{c}_{i}\}$ is the parameter for the $i$-th Gaussian. To render a set of 3D Gaussians, we need to project them into the image plane as 2D Gaussians. Volumetric rendering is then performed for each pixel in front-to-back depth order to evaluate the final color and alpha. In this work, we use the highly optimized renderer implementation from Kerbl et al. (2023) to optimize ${\Theta}$.
We initialize the 3D Gaussians with random positions sampled inside a sphere, with unit scaling and no rotation. These 3D Gaussians are periodically densified during optimization. Different from the reconstruction pipeline, we start from fewer Gaussians but densify it more frequently to align with the generation progress. We follow the recommended practices from previous works (Poole et al., 2022; Huang et al., 2023; Lin et al., 2023) and use SDS to optimize the 3D Gaussians. At each step, we sample a random camera pose $p$ orbiting the object center, and render the RGB image $I^{p}_{\text{RGB}}$ and transparency $I^{p}_{\text{A}}$ of the current view. Similar to Dreamtime (Huang et al., 2023), we decrease the timestep $t$ linearly, which is used to weight the random noise $\epsilon$ added to the rendered RGB image. Then, different 2D diffusion priors $\phi$ are used to guide the SDS denoising step, which is back-propagated to the 3D Gaussians.
Image-to-3D. For the image-to-3D task, an image $\tilde{I}^{r}_{\text{RGB}}$ and a foreground mask $\tilde{I}^{r}_{\text{A}}$ are given as input. Zero-1-to-3 XL (Liu et al., 2023b; Deitke et al., 2023b) is adopted as the 2D diffusion prior. The SDS loss can be formulated as:
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,p,\mathbf{\epsilon}}\left[(\epsilon_{\phi}(I^{p}_{\text{RGB}};t,\tilde{I}^{r}_{\text{RGB}},\Delta p)-\epsilon)\frac{\partial I^{p}_{\text{RGB}}}{\partial{\Theta}}\right]$ (1)
where $\epsilon_{\phi}(\cdot)$ is the predicted noise by the 2D diffusion prior $\phi$, and $\Delta p$ is the relative camera pose change from the reference camera $r$. Additionally, we optimize the reference view image $I_{\text{RGB}}^{r}$ and transparency $I_{\text{A}}^{r}$ to align with the input:
$\mathcal{L}_{\text{Ref}}=\lambda_{\text{RGB}}||I_{\text{RGB}}^{r}-\tilde{I}_{\text{RGB}}^{r}||^{2}_{2}+\lambda_{\text{A}}||I_{\text{A}}^{r}-\tilde{I}_{\text{A}}^{r}||^{2}_{2}$ (2)
where $\lambda_{\text{RGB}}$ and $\lambda_{\text{A}}$ are the weights which are linearly increased during training. The final loss is the weighted sum of the above three losses.
Text-to-3D. The input for text-to-3D is a single text prompt. Following previous works, Stable-diffusion (Rombach et al., 2022) is used for the text-to-3D task. The SDS loss can be formulated as:
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,p,\mathbf{\epsilon}}\left[(\epsilon_{\phi}(I^{p}_{\text{RGB}};t,e)-\epsilon)\frac{\partial I^{p}_{\text{RGB}}}{\partial{\Theta}}\right]$ (3)
where $e$ is the CLIP embeddings of the input text description.
Dissusion. We observe that the generated Gaussians often look blurry and lack details even with longer SDS training iterations. This could be explained by the ambiguity of SDS loss. Since each optimization step may provide inconsistent 3D guidance, it’s hard for the algorithm to correctly densify the under-reconstruction regions or prune over-reconstruction regions as in reconstruction. This observation leads us to the following mesh extraction and texture refinement designs.
3.2 Efficient Mesh Extraction
Polygonal mesh is a widely used 3D representation, particularly in industrial applications. Many previous works (Poole et al., 2022; Lin et al., 2023; Tsalicoglou et al., 2023; Tang et al., 2023a) export the NeRF representation into a mesh-based representation for high-resolution fine-tuning. We also seek to convert the generated 3D Gaussians into meshes and further refine the texture.
To the best of our knowledge, the polygonal mesh extraction from 3D Gaussians is still an unexplored problem. Since the spatial density is described by a large number of 3D Gaussians, brute-force querying of a dense 3D density grid can be slow and inefficient. It’s also unclear how to extract the appearance in 3D, as the color blending is only defined with projected 2D Gaussians (Kerbl et al., 2023). Here, we propose an efficient algorithm to extract a textured mesh based on block-wise local density query and back-projected color.
Local Density Query. To extract the mesh geometry, a dense density grid is needed to apply the Marching Cubes (Lorensen & Cline, 1998) algorithm. An important feature of the Gaussian splatting algorithm is that over-sized Gaussians will be split or pruned during optimization. This is the foundation of the tile-based culling technique for efficient rasterization (Kerbl et al., 2023). We also leverage this feature to perform block-wise density queries.
We first divide the 3D space of $(-1,1)^{3}$ into $16^{3}$ blocks, then cull the Gaussians whose centers are located outside each local block. This effectively reduces the total number of Gaussians to query in each block. We then query a $8^{3}$ dense grid inside each block, which leads to a final $128^{3}$ dense grid. For each query at grid position $\mathbf{x}$, we sum up the weighted opacity of each remained 3D Gaussian:
$d(\mathbf{x})=\sum_{i}\alpha_{i}\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{x_{i}})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{x_{i}}))$ (4)
where $\Sigma_{i}$ is the covariance matrix built from scaling $\mathbf{s}_{i}$ and rotation $\mathbf{q}_{i}$. An empirical threshold is then used to extract the mesh surface through Marching Cubes. Decimation and Remeshing (Cignoni et al., 2008) are applied to post-process the extracted mesh to make it smooth.
Color Back-projection. Since we have acquired the mesh geometry, we can back project the rendered RGB image to the mesh surface and bake it as the texture. We first unwrap the mesh’s UV coordinates (Young, 2021) and initialize an empty texture image. Then we uniformly choose 8 azimuths and 3 elevations, plus the top and bottom views to render the corresponding RGB image. Each pixel from these RGB images can be back-projected to the texture image based on the UV coordinate. Following Richardson et al. (2023), we exclude the pixels with a small camera space z-direction normal to avoid unstable projection at mesh boundaries. This back-projected texture image serves as an initialization for the next mesh texture fine-tuning stage.
Figure 3: Different Texture Fine-tuning Objectives. We show that SDS loss produces artifacts for UV space texture optimization, while the proposed MSE loss avoids this.
3.3 UV-space Texture Refinement
Due to the ambiguity of SDS optimization, the meshes extracted from 3D Gaussians usually possess blurry texture as illustrated in Figure 3. We therefore propose a second stage to refine the texture image. However, fine-tuning the UV-space directly with SDS loss often leads to artifacts, which is also observed in previous works (Liao et al., 2023). This is due to the mipmap texture sampling technique used in differentiable rasterization (Laine et al., 2020). With ambiguous guidance like SDS, the gradient propagated to each mipmap level results in over-saturated color blocks. Therefore, we seek more definite guidance to fine-tune a blurry texture.
We draw inspiration from the image-to-image synthesis of SDEdit (Meng et al., 2021) and the reconstruction settings. Since we already have an initialization texture, we can render a blurry image $I^{p}_{\text{coarse}}$ from an arbitrary camera view $p$. Then, we perturb the image with random noise and apply a multi-step denoising process $f_{\phi}(\cdot)$ using the 2D diffusion prior to obtaining a refined image:
$I^{p}_{\text{fine}}=f_{\phi}(I^{p}_{\text{coarse}}+\epsilon(t_{\text{start}});t_{\text{start}},c)$ (5)
where $\epsilon(t_{\text{start}})$ is a random noise at timestep $t_{\text{start}}$, $c$ is $\Delta p$ for image-to-3D and $e$ for text-to-3D respectively. The starting timestep $t_{\text{start}}$ is carefully chosen to limit the noise strength, so the refined image can enhance details without breaking the original content. This refined image is then used to optimize the texture through a pixel-wise MSE loss:
$\mathcal{L}_{\text{MSE}}=||I^{p}_{\text{fine}}-I^{p}_{\text{coarse}}||^{2}_{2}$ (6)
For image-to-3D tasks, we still apply the reference view RGBA loss in Equation 2. We find that only about 50 steps can lead to good details for most cases, while more iterations can further enhance the details of the texture.
4 Experiments
4.1 Implementation Details
We train $500$ steps for the first stage and $50$ steps for the second stage. The 3D Gaussians are initialized to $0.1$ opacity and grey color inside a sphere of radius $0.5$. The rendering resolution is increased from $64$ to $512$ for Gaussian splatting, and randomly sampled from $128$ to $1024$ for mesh. The loss weights for RGB and transperency are linearly increased from 00 to $10^{4}$ and $10^{3}$ during training. We sample random camera poses at a fixed radius of $2$ for image-to-3D and $2.5$ for text-to-3D, y-axis FOV of $49$ degree, with the azimuth in $[-180,180]$ degree and elevation in $[-30,30]$ degree. The background is rendered randomly as white or black for Gaussian splatting. For image-to-3D task, the two stages each take around 1 minute. We preprocess the input image by background removal (Qin et al., 2020) and recenter the foreground object. The 3D Gaussians are initialized with $5000$ random particles and densified for each $100$ steps. For text-to-3D task, due to the larger resolution of $512\times 512$ used by Stable Diffusion (Rombach et al., 2022) model, each stage takes around 2 minutes to finish. We initialize the 3D Gaussians with $1000$ random particles and densify them for each $50$ steps. For mesh extraction, we use an empirical threshold of $1$ for Marching Cubes. All experiments are performed and measured with an NVIDIA V100 (16GB) GPU, while our method requires less than 8 GB GPU memory. Please check the supplementary materials for more details.
Figure 4: Comparisons on Image-to-3D. Our method achieves a better balance between generation speed and mesh quality on various images.
Figure 5: Comparisons on Text-to-3D. For DreamFusion, we use the implementation from Guo et al. (2023) which also uses Stable-Diffusion as the 2D prior.
Type CLIP-Similarity $\uparrow$ Generation Time $\downarrow$
One-2-3-45 (Liu et al., 2023a) Inference-only 0.594 45 seconds
Point-E (Nichol et al., 2022) Inference-only 0.587 78 seconds
Shap-E (Jun & Nichol, 2023) Inference-only 0.591 27 seconds
Zero-1-to-3 (Liu et al., 2023b) Optimization-based 0.647 20 minutes
$*{}^{*}$ (Liu et al., 2023b) Optimization-based 0.778 30 minutes
Ours (Stage 1 Only) Optimization-based 0.678 1 minute
Ours Optimization-based 0.738 2 minutes
Table 1: Quantitative Comparisons on generation quality and speed for image-to-3D tasks. For $*{}^{*}$, a mesh fine-tuning stage is used to further improve quality (Tang, 2022).
4.2 Qualitative Comparisons
We first provide qualitative comparisons on image-to-3D in Figure 4. We primarily compare with three baselines from both optimization-based methods (Liu et al., 2023b) and inference-only methods (Liu et al., 2023a; Jun & Nichol, 2023). For all compared methods, we export the generated models as polygonal meshes with vertex color or texture images, and render them under ambient lighting. In terms of generation speed, our approach exhibits a noteworthy acceleration when contrasted with optimization-based methods. Regarding the quality of generated models, our method outperforms inference-only methods especially with respect to the fidelity of 3D geometry and visual appearance. In general, our method achieves a better balance between generation quality and speed, reaching comparable quality as optimization-based methods while only marginally slower than inference-only methods. In Figure 5, we compare the results on text-to-3D. Consistent with our findings in image-to-3D tasks, our method achieves better quality than inference-based methods and faster speed than other optimization-based methods. Furthermore, we highlight the quality of our exported meshes in Figure 6. These meshes exhibit uniform triangulation, smooth surface normals, and clear texture images, rendering them well-suited for seamless integration into downstream applications. For instance, leveraging software such as Blender (Community, 2018), we can readily employ these meshes for rigging and animation purposes.
4.3 Quantitative Comparisons
In Table 1, we report the CLIP-similarity (Radford et al., 2021; Qian et al., 2023; Liu et al., 2023a) and average generation time of different image-to-3D methods on a collection of images from previous works (Melas-Kyriazi et al., 2023; Liu et al., 2023a; Tang et al., 2023b) and Internet. We also conduct an user study on the generation quality detailed in Table 2. This study centers on the assessment of reference view consistency and overall generation quality, which are two critical aspects in the context of image-to-3D tasks. Our two-stage results achieve better view consistency and generation quality compared to inference-only methods. Although our mesh quality falls slightly behind that of other optimization-based methods, we reach a significant acceleration of over 10 times.
Figure 6: Mesh Exportation. We export high quality textured mesh from 3D Gaussians, which can be seamlessly used in downstream applications like rigged animation.
Figure 7: Ablation Study. We ablate the design choices in stage 1 training.
Zero-1-to-3 One-2-3-45 Shap-E Ours
Ref. View Consistency $\uparrow$ 3.43 2.33 1.81 4.25
Overall Model Quality $\uparrow$ 3.02 1.86 1.52 3.84
Table 2: User Study on image-to-3D tasks. The rating is of scale 1-5, the higher the better.
4.4 Ablation Study
We carry out ablation studies on the design of our methods in Figure 7. We are mainly interested in the generative Gaussian splatting training, given that mesh fine-tuning has been well explored in previous methods (Tang et al., 2023a; Lin et al., 2023). Specifically, we perform ablation on three aspects of our method: 1) Periodical densification of 3D Gaussians. 2) Linear annealing of timestep $t$ for SDS loss. 3) Effect of the reference view loss $\mathcal{L}_{\text{Ref}}$. Our findings reveal that the omission of any of these design elements results in a degradation of the quality of the models generated in the first stage. Specifically, the final Gaussians exhibit increased blurriness and inaccuracies, which further affects the second fine-tuning stage.
5 Limitations and Conclusion
In this work, we present DreamGausssion, a 3D content generation framework that significantly improves the efficiency of 3D content creation. The key features of our work are: 1) We design the generative Gaussian splatting pipeline which is highly efficient for 3D generation. 2) We propose a mesh extraction algorithm that effectively derives textured meshes from 3D Gaussians. 3) With our texture fine-tuning stage, our work can produce ready-to-use 3D assets with high-quality polygonal meshes from either a single image or text description within a few minutes.
Limitations. We share common problems with previous text-to-3D works: Multi-face Janus problem and baked lighting. Fortunately, it’s possible to address these problems with recent advances in multi-view 2D diffusion models (Shi et al., 2023; Liu et al., 2023c; Zhao et al., 2023) and latent BRDF auto-encoder (Xu et al., 2023b). Besides, the back-view texture generated in our image-to-3D results may look blurry. This can be alleviated with longer stage 2 training.
Acknowledgments
This work is supported by the Sichuan Science and Technology Program (2023YFSY0008), National Natural Science Foundation of China (61632003, 61375022, 61403005), Grant SCITLAB-20017 of Intelligent Terminal Key Laboratory of SiChuan Province, Beijing Advanced Innovation Center for Intelligent Robots and Systems (2018IRS11), and PEK-SenseTime Joint Laboratory of Machine Vision. This project is also funded by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix A Appendix
A.1 More Implementation Details
Learning Rate. For the learning rate of Gaussian splatting, we set different values for different parameters. The learning rate for position is decayed from $1\times 10^{-3}$ to $2\times 10^{-5}$ in $500$ steps, for feature is set to $0.01$, for opacity is $0.05$, for scaling and rotation is $5\times 10^{-3}$. For mesh texture fine-tuning, the learning rate for texture image is set to $0.2$. We use the Adam (Kingma & Ba, 2014) optimizer for both stages.
Densification and Pruning. Following  Kerbl et al. (2023), the densification in image-to-3D is applied for Gaussians with accumulated gradient larger than $0.5$ and max scaling smaller than $0.05$. In text-to-3D, we set the gradient threshold to $0.01$ to encourage densification. We also prune the Gaussians with an opacity less than $0.01$ or max scaling larger than $0.05$.
Evaluation Settings. We adopt the CLIP-similarity metric (Melas-Kyriazi et al., 2023; Qian et al., 2023; Liu et al., 2023a) to evaluate the image-to-3D quality. A dataset of 30 images collected from previous works (Melas-Kyriazi et al., 2023; Liu et al., 2023a; Tang et al., 2023b; Liu et al., 2023c) and Internet covering various objects is used. We then render $8$ views with uniformly sampled azimuth angles $[0,45,90,135,180,225,270,315]$ and zero elevation angle. These rendered images are used to calculate the CLIP similarities with the reference view, and we average the different views for the final metric. We use the laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [^1] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k checkpoint to calculate CLIP similarity. For the user study, we render 360 degree rotating videos of 3D models generated from a collection of 15 images. There are in total 60 videos for 4 methods (Zero-1-to-3 (Liu et al., 2023b), One-2-3-45 Liu et al. (2023a), Shap-E Jun & Nichol (2023), and our method) to evaluate. Each volunteer is shown 15 samples containing the input image and a rendered video from a random method, and ask them to rate in two aspects: reference view consistency and overall model quality. We collect results from 40 volunteers and get 600 valid scores in total.
A.2 More Results
In Figure 8, we show more visualization results of our method. Specially, we compare the mesh output before and after our texture fine-tuning stage. We also compare against a SDS-based mesh fine-tuning method for Zero-1-to-3 (Liu et al., 2023b) noted as $*{}^{*}$ (Tang, 2022). Both stages of our method are faster than previous two-stage image-to-3D methods, while still reaching comparable generation quality.
Our method also support images with non-zero elevations. As illustrated in Figure 10, our method can perform image-to-3D correctly with an extra estimated elevation angle as input. We make sure the random elevation sampling covers the input elevation and at least $[-30,30]$ degree.
In Figure 11, we demostrate the text-to-image-to-3D pipeline (Liu et al., 2023a; Qian et al., 2023). We first apply text-to-image diffusion models (Rombach et al., 2022) to synthesize an image given a text prompt, then perform image-to-3D using our model. This usually gives better results compared to directly performing text-to-3D pipeline, and takes less time to generate. We show more animation results from our exported meshes in Figure 9.
We also illustrate the limitations of our method in Figure 12. Our image-to-3D pipeline may produce blurry back-view image and cannot generate fine details, which looks unmatched to the front reference view. With longer training of stage 2, the blurry problem of back view can be alleviated. For text-to-3D, we share common problems with previous methods, including the multi-face Janus problem and baked lighting in texture images.
Figure 8: More Qualitative Comparisons. We compare the results from two training stages of our method and Zero-1-to-3 (Liu et al., 2023b).
Figure 9: Results on mesh animation. Our exported meshes are ready-to-use for downstream applications like rigged animation.
Figure 10: Results on images with different elevations. Our method supports input images with a non-zero elevation angle.
Figure 11: Text-to-image-to-3D. We first synthesize an image given a text prompt, then perform image-to-3D generation.
Figure 12: Limitations. Visualization of the limitations of our method.